OpenStudy (anonymous):

Assuming that ∫[0,1,,] f(x)dx=3 ∫[0,2] f(x)dx=5 ∫[1,4] f(x)dx=8 What is ∫[4,2] f(x)dx?

OpenStudy (unklerhaukus):

\[\int\limits_4^2=-\int\limits_1^4-\int\limits_0^1+\int\limits_0^2\]